# Non-crossing permutations for the KP solitons under the Gel’fand-Dickey reductions and the vertex operators

### v1

@@ 4.1. The Boussinesq equation from the KP theory
[…]
@@
(4.8) {$&x~{}\to~{}x+c_{1}t_{3},\\ &t_{2}~{}\to~{}t_{2}+c_{2}t_{3},\\ &t_{3}~{}\to~{}{t_{3}},\end{aligned}\end{array}\right.\qquad\text{which gives}\qquad\left\{\begin{array}[]{ll}\vspace{1ex}\begin{aligned} &\frac{\partial}{\partial x}~{}\to~{}\frac{\partial}{\partial{x}},\quad\quad\frac{\partial}{\partial t_{2}}~{}\to~{}\frac{\partial}{\partial{t_{2}}},\\ &\frac{\partial}{\partial t_{3}}~{}\to~{}\frac{\partial}{\partial t_{3}}+c_{1}\frac{\partial}{\partial{x}}+c_{2}\frac{\partial}{\partial{t_{2}}},\end{aligned}\end{array}\right.$
@@ 4.2. One soliton solution of the Boussinesq equation
    $-\frac{1}{2}c_{0}<v_{[1,3]}<\frac{1}{2}c_{0},\qquad\frac{9}{8}c_{0}^{2}<a_{[1,3]}\leq\frac{3}{2}c_{0}^{2},$
@@
    $-c_{0}<v_{[2,3]}<-\frac{1}{2}c_{0},\qquad 0<a_{[2,3]}<\frac{9}{8}c_{0}^{2}.$
Note that the amplitude $a_{[i,j]}$ of the soliton is limited by $23c_{0}^{2}/$ 2.
@@ 4.4. Discussions
In this section, we studied the (good) Boussinesq equation and classified the general solutions of the equation. We hope that our results provide a model of bidirectional soliton gas as discussed recently in [1]. We also found that there exists at most one resonant solution (Y-shape soliton) or one slow propagating soliton with large amplitude for the regular solitons (see [13]). It might be interesting to discuss the effect of those special solitons among two groups of counter propagating solitons. However, it may not be so physical if the (good) Boussinesq equation is used for a shallow water wave model. This is because that a larger soliton has a slower velocity, and even that the largest soliton has the zero velocity.
@@ 5. Vertex operator construction of the KP solitons under the $\ell$-reductions
$X(p_{i},q_{j})=e^{\varphi(\mathbf{t},p_{i},q_{j})}e^{-\varphi(\tilde{\partial},p_{i}^{-1},q_{j}^{-1})},\quad\text{where}\quad\varphi(\mathbf{t},p_{i},q_{j}):=\mathop{\textstyle\sum}\limits_{n=1}^{\infty}(q_{j}^{n}-p_{i}^{n})t_{n}.$
The following lemma is well-known and easy to show.
@@
where the normal ordering symbol ::::$\cdot$:: ⋅ : implies to move differential operators $\partial_{n}$ to the right. It then follows the following properties
@@ 5.1. The vertex operators on $\text{Gr}(n,m)_{\geq 0}$
In the present paper, the parameters $\{p_{i},q_{j}\}$ are determined by the roots of the spectral curve $\Phi_{\ell}(\kappa,\alpha)=0$ for each $\alpha$, i.e. $p_{i}=\kappa_{i^{\prime}}[\alpha]$ and $q_{j}=\kappa_{j^{\prime}}[\alpha]$ for some $i^{\prime},j^{\prime}\in\mathcal{I}[\alpha]$. As in the previous section, we take $2\leq|\mathcal{I}[\alpha]|=m\leq\ell$, and consider $11\leq n\leq m-$ 1. We take an (irreducible) element $A[\alpha]\in\text{Gr}(n,m)_{\geq 0}$. Let $\mathcal{I}_{P}[\alpha]$ and $\mathcal{I}_{Q}[\alpha]$ be the sets of pivots and non-pivots of the matrix $A[\alpha]$, i.e. $\mathcal{I}[\alpha]=\mathcal{I}_{P}[\alpha]\sqcup\mathcal{I}_{Q}[\alpha]$. We use the following labels for the nonzero elements in $A[\alpha]$,
$\mathcal{I}_{P}[\alpha]$ =$=\{i_{1}<i_{2}<\cdots<i_{n}\},$
@@
where $j_{l}^{(k)}$ is the column index of the nonzero element in the $k$-th row and $j_{l}$-th column. That is, the nonzero elements except the pivots in the matrix $A[\alpha]$ are
[…]
@@
Note here that the parameters $(p_{i},q_{j})$ in $X(p_{i},q_{j})$ are from $(i\in\mathcal{I}_{P}[\alpha],j\in\mathcal{I}_{Q}[\alpha])$, so that these $X(p_{i},q_{j})$ satisfy the condition in Lemma 5.1. Then we have the following proposition.
@@
Proof. We first note (Theorem 3.10 in [8]) that we have $\tau(\mathbf{t})/E_{I_{0}}(\mathbf{t})=\tilde{\vartheta}_{\tilde{g}}(z;\tilde{\Omega})$, called the $M$-theta function defined by
$\tilde{\vartheta}_{\tilde{g}}(z;\tilde{\Omega})$ $=\mathop{\textstyle\sum}\limits_{{\bf m}\in\{0,1\}^{\tilde{g}}}\exp 2\pi i\left(\mathop{\textstyle\sum}\limits_{j<k}^{\tilde{g}}m_{j}m_{k}\tilde{\Omega}_{j,k}+\mathop{\textstyle\sum}\limits_{j=1}^{\tilde{g}}m_{j}z_{j}\right)$
@@
where we have used the order (5.1). Then for $z_{j}$ with $j=\tilde{g}_{k-1}+l$, the $e^{2\pi iz_{j}}$ in $\tilde{\vartheta}_{\tilde{g}}(z,\tilde{\Omega})$ becomes
@@ 5.2. The vertex operator construction of the KP solitons under the $\ell$-reduction
$\tau(\mathbf{t},\alpha_{1},\alpha_{2})=V[\alpha_{1}]\cdot V[\alpha_{2}]\cdot 1.$
@@
Notice that both $b_{i,j}$ and $c_{i,j}$ are positive. This difference implies that they have different phase shifts for solitons contained in the solutions due to the different values of the free parameters. We might say that they are topologically the same, and we say $V[\alpha_{1},\alpha_{2}]\equiv V[\alpha_{1}]\cdot V[\alpha_{2}]$.
Acknowledgements. The authors appreciate a research fund from Shandong University of Science and Technology. One of the authors (C.L) is supported by National Natural Science Foundation of China (Grant No. 12071237).

### v2

@@ 4.1. The Boussinesq equation from the KP theory
(4.2) $\left\{\begin{array}[]{ll}\vskip 4.30554pt\begin{aligned} &\frac{\partial u}{\partial t_{2}}=\frac{4}{3}w_{x},\\ &\frac{\partial w}{\partial t_{2}}=-\frac{1}{4}u_{xxx}-\frac{3}{2}uu_{x}.\end{aligned}\end{array}\right$.
@@
(4.8) $\left\{\begin{array}[]{ll}\vskip 4.30554pt\begin{aligned} &x~\to~x+c_{1}t_{3},\\ &t_{2}~\to~t_{2}+c_{2}t_{3},\\ &t_{3}~\to~{t_{3}},\end{aligned}\end{array}\right.\qquad\text{which gives}\qquad\left\{\begin{array}[]{ll}\vskip 4.30554pt\begin{aligned} &\frac{\partial}{\partial x}~\to~\frac{\partial}{\partial{x}},\quad\quad\frac{\partial}{\partial t_{2}}~\to~\frac{\partial}{\partial{t_{2}}},\\ &\frac{\partial}{\partial t_{3}}~\to~\frac{\partial}{\partial t_{3}}+c_{1}\frac{\partial}{\partial{x}}+c_{2}\frac{\partial}{\partial{t_{2}}},\end{aligned}\end{array}\right$.
@@ 4.2. One soliton solution of the Boussinesq equation
    $\frac{1}{2}c_{0}<v_{[1,3]}<\frac{1}{2}c_{0},\qquad\frac{9}{8}c_{0}^{2}<a_{[1,3]}\leq\frac{3}{2}c_{0}^{2},$
@@
    $c_{0}<v_{[2,3]}<-\frac{1}{2}c_{0},\qquad 0<a_{[2,3]}<\frac{9}{8}c_{0}^{2}.$
Note that the amplitude $a_{[i,j]}$ of the soliton is limited by $3c_{0}^{2}/2$.
@@ 4.4. Discussions
In this section, we studied the (good) Boussinesq equation and classified the general solutions of the equation. We hope that our results provide a model of bidirectional soliton gas as discussed recently in [4]. We also found that there exists at most one resonant solution (Y-shape soliton) or one slow propagating soliton with large amplitude for the regular solitons (see [16]). It might be interesting to discuss the effect of those special solitons among two groups of counter propagating solitons. However, it may not be so physical if the (good) Boussinesq equation is used for a shallow water wave model. This is because that a larger soliton has a slower velocity, and even that the largest soliton has the zero velocity. It was shown in [2] that these slow solitons are unstable under a small perturbation (also see [3]).
@@ 5. Vertex operator construction of the KP solitons under the $\ell$-reductions
$X(p_{i},q_{j})=e^{\varphi(\mathbf{t},p_{i},q_{j})}e^{-\varphi(\tilde{\partial},p_{i}^{-1},q_{j}^{-1})},\quad\text{where}\quad\varphi(\mathbf{t},p_{i},q_{j}):=\mathop{\textstyle\sum}\limits_{n=1}^{\infty}(q_{j}^{n}-p_{i}^{n})t_{n},$
where $\tilde{\partial}=(\partial_{x},\frac{1}{2}\partial_{t_{2}},\cdots)$. The following lemma is well-known and easy to show.
@@
where the normal ordering symbol $\cdot:$ implies to move differential operators $\partial_{n}$ to the right. It then follows the following properties
@@ 5.1. The vertex operators on $\text{Gr}(n,m)_{\geq 0}$
In the present paper, the parameters $\{p_{i},q_{j}\}$ are determined by the roots of the spectral curve $\Phi_{\ell}(\kappa,\alpha)=0$ for each $\alpha$, i.e. $p_{i}=\kappa_{i^{\prime}}[\alpha]$ and $q_{j}=\kappa_{j^{\prime}}[\alpha]$ for some $i^{\prime},j^{\prime}\in\mathcal{I}[\alpha]$. As in the previous section, we take $2\leq|\mathcal{I}[\alpha]|=m\leq\ell$, and consider $1\leq n\leq m-1$. We take an (irreducible) element $A[\alpha]\in\text{Gr}(n,m)_{\geq 0}$. Let $\mathcal{I}_{P}[\alpha]$ and $\mathcal{I}_{Q}[\alpha]$ be the sets of pivots and non-pivots of the matrix $A[\alpha]$, i.e. $\mathcal{I}[\alpha]=\mathcal{I}_{P}[\alpha]\sqcup\mathcal{I}_{Q}[\alpha]$. We use the following labels for the nonzero elements in $A[\alpha]$,
$\mathcal{I}_{P}[\alpha]$ $=\{i_{1}<i_{2}<\cdots<i_{n}\},$
@@
where $j_{l}^{(k)}$ is the column index of the nonzero element in the $k$-th row and $j_{l}^{(k)}$-th column. That is, the nonzero elements except the pivots in the matrix $A[\alpha]$ are
$\left\{\,a_{k,j^{(k)}_{l}}:\,1\leq l\leq q_{k},~1\leq k\leq n\,\right\}.$
@@
Note here that the parameters $(\kappa_{i},\kappa_{j})$ in $X(\kappa_{i},\kappa_{j})$ are from $(i\in\mathcal{I}_{P}[\alpha],j\in\mathcal{I}_{Q}[\alpha])$, so that these $X(\kappa_{i},\kappa_{j})$ satisfy the condition in Lemma 5.1. Then we have the following proposition.
@@
[…]
@@
$\phi_{\tilde{g}_{k-1}+l}(\mathbf{t})=\varphi(\mathbf{t},\kappa_{i_{k}},\kappa_{j_{l}^{(k)}}),\quad\phi^{0}_{\tilde{g}_{k-1}+l}=\ln\left(b_{k,j_{l}^{(k)}}\right),\qquad(1\leq l\leq q_{k},~\,1\leq k\leq n),$
@@
[…]
@@ 5.2. The vertex operator construction of the KP solitons under the $\ell$-reduction
$\tau(\mathbf{t};\alpha_{1},\alpha_{2})=V[\alpha_{1}]\cdot V[\alpha_{2}]\cdot 1.$
@@
Notice that both $b_{i,j}$ and $c_{i,j}$ are positive. The solutions generated by $V[\alpha_{1}]\cdot V[\alpha_{2}]$ and $V[\alpha_{1},\alpha_{2}]$ are both regular and have the same set of solitons. However, the phases of the solitons are different due to the different positive matrices in the vertex operators. We might say that they are topologically the same, and we may say $V[\alpha_{1},\alpha_{2}]\equiv V[\alpha_{1}]\cdot V[\alpha_{2}]$.
Acknowledgements. We would like to thank the referees for valuable comments and informing us the references [1] and [2]. This article is dedicated to Professor Ke Wu at Capital Normal University in celebration of his 80th birthday. One of the authors (C.L.) is supported by National Natural Science Foundation of China (Grant No. 12071237).